2mL of 5(M) H3PO4 is mixed with 3mL 2(M) Ca(OH)2, what will be the strength of the resulting solution?

Grade:12

1 Answers

Vikas TU
6 years ago
As the moles wpuld be constant after and beforemixing of the sol. both,
M1V1 = M2V2
Therefore,
2*5 + 3*2 = M*5
16 = 5M
M = 16/5 => 3.2M
This would be the final sstrength of the soln.